### v1

@@ Abstract
Beam squint poses a fundamental challenge in wideband hybrid beamforming, particularly for mmWave and THz systems that demand both ultra-wide bandwidth and high directional beams. While conventional phase shifter-based beamformers may offer partial mitigation, True Time Delay (TTD) units provide a fundamentally more effective solution by enabling frequency-independent beam steering. However, the high cost of TTD units has recently driven much interest in Sparse-TTD architectures, which combine a limited number of TTDs with a higher number of conventional phase shifters to balance performance and cost. This paper provides a critical examination of beam squint mitigation strategies in wideband hybrid beamformers, comparing Full-TTD, Sparse-TTD, and Non-TTD architectures. We analyze recent Non-TTD approaches, specifically the scheme leveraging the wideband beam gain (WBBG) concept, evaluating their performance and cost characteristics against TTD-based solutions. A key focus is placed on the practical limitations of Sparse-TTD architectures, particularly the often-overlooked requirement for wideband phase shifters operating alongside TTDs, which can significantly impact performance and implementation cost in real-world scenarios, especially for ultra-wideband applications. Finally, we conduct a cost-performance analysis to examine the trade-offs inherent in each architecture and provide guidance on selecting the most suitable hybrid beamforming structure for various fractional bandwidth regimes.
Index Terms:
Beam squint, hybrid beamforming, True Time Delay (TTD), wideband phase shifter, wideband beam gain.
@@ I Introduction
Efficient beamforming is essential for enabling 6G technologies that exploit mmWave and THz bands to deliver ultra-high data rates and dense connectivity.In phased array systems, analog beamforming adjusts signal phases of antenna elements utilizing a single RF chain, while digital beamforming offers fine-grained amplitude and phase control at the cost of one RF chain per antenna element. To balance performance and hardware complexity, hybrid beamforming combines both approaches by connecting fewer RF chains to large antenna arrays via analog phase shifters. This allows flexible beam control with reduced cost and power consumption. In contrast to conventional beamforming methods, which often rely on narrowband assumptions, contemporary applications increasingly demand wideband and ultra-wideband beamforming to support ultra-high data rates as well as robust communications. For instance, in an ultra-wideband multi-carrier transmitter, frequency diversity enhances robustness by allowing the system to maintain performance despite deep fading or interference on specific carriers, as unaffected carriers can ensure reliable operation [1]. This is particularly vital for frequency-selective fading channels, driving significant research into advanced wideband beamforming techniques that underpin the evolution of resilient, high-capacity 6G systems. Despite its advantages, wideband beamforming introduces a critical challenge known as beam squint. Beam squint arises when fixed phase shifts optimized for a target frequency $f_{0}$, lead to frequency-dependent beam direction change for frequency $f\neq f_{0}$. For the simple case of a uniform linear array (ULA) transmitter, the main lobe of the radiated beam at frequency $f$ is deviated by $\Delta\theta=\sin^{-1}(\frac{f_{0}}{f}\sin(\theta_{0}))-\theta_{0}$, where $\theta_{0}$ corresponds to the intended beam angle at frequency $f_{0}$ with respect to the boresight of the antenna. This frequency-dependent beam squint degrades beamforming gain and overall system performance, particularly in ultra-wideband systems.
@@
To mitigate beam squint in wideband radar and communication systems, several techniques have been proposed. The first solution scheme focuses on implementing signal processing and beamforming techniques implemented using conventional analog-domain phase shifters. For instance, in the context of MIMO wideband radar, the work in [2] addresses beam pattern design under a space-frequency nulling constraint, aiming to suppress frequency-dependent interference caused by beam squint. In [3], the authors propose a beamforming scheme for wideband radars that synthesizes frequency-invariant beam patterns via constrained waveform design, incorporating nulling and peak-to-average power ratio control in the space-frequency domain. More recently, [4] introduced a hybrid beamforming approach that broadens the beam in the angular domain to achieve max-min fair beam gain across spatially squinted beams.
@@
Wideband beamformers using conventional phase shifters can partially mitigate beam squint, as they cannot provide an optimal solution due to the group delay associated with such components. In this regard, the second approach aims to more fundamentally mitigate beam squint by replacing phase shifters with true time delays (TTDs). Unlike phase shifters that introduce frequency-dependent phase shifts, TTDs apply frequency-independent delays, ensuring consistent beam direction across a wide bandwidth. In general, the beam squint mitigation through Non-TTD-based hybrid structures such as the one presented in [4] typically follows the general block diagram of hybrid beamforer illustrated in Fig. 1-a. In contrast, TTD-based hybrid architectures may adopt a Full-TTD structure with no phase shifters, as shown in Fig. 1-b [5], or employ Sparse-TTD structures in which the number of TTDs is smaller than the number of phase shifters. Examples of Sparse-TTD configurations include those shown in Fig. 1-c [6, 7], Fig. 1-d [8], and Fig. 1-e [9].
@@
Given the relatively high cost of TTDs compared to phase shifters, a Sparse-TTD configuration aims to reduce the hardware cost of a Full-TTD wideband hybrid beamformer. However, this approach presents several notable drawbacks. First, it does not completely eliminate beam squint, leaving residual misalignment across some frequency bands, since full squint mitigation requires equipping each antenna element with a dedicated TTD. Second, the joint optimization of three sets of variables, namely the baseband precoders, the delay parameters of the TTDs, and the tuning parameters of the phase shifters, introduces considerable computational complexity, as the design must jointly optimize all three component vectors to achieve near-optimal performance. Finally, and most critically, a challenge often overlooked in the literature is the requirement for wideband phase shifters to operate alongside TTDs. This necessity introduces additional implementation concerns, which will be examined in detail in this work. In light of the above, we aim to reassess the practicality of existing wideband hybrid beamforming architectures under real-world design constraints. Our analysis takes into account a range of physical parameters, implementation challenges, and the trade-offs between cost, complexity, and performance.
The main contributions of this paper are summarized as follows:
@@
  • While Sparse-TTD wideband hybrid beamformers are widely regarded in the literature as a high-performance cost-effective solution for beam squint mitigation, we critically examine their practical limitations. In particular, we highlight the often-overlooked requirement for wideband phase shifters and discuss how this constraint can significantly affect performance and implementation cost in real-world scenarios. These are explored in Sections II and III.
@@
Figure 1: Wideband hybrid beamformer leveraging (a) conventional phase-shifters (PSs) without deploying TTDs (Non-TTD), (b) one TTD per antenna element (Full-TTD), and (c): sparse TTDs together with a block of phase-shifters (Sparse-TTD)
@@ II Sparse-TTD Beamforming: Implementation Challenges
$\max_{\mathbf{D}_{m},\mathbf{T},\mathbf{A}}$ $Q(\mathbf{D}_{m},\mathbf{T},\mathbf{A})$
@@
where $\mathbf{D}_{m}$ represents the digital beamformer corresponding to each carrier $m$, $\mathbf{T}$ is the vector encompassing the tunable delays of the TTDs ($0\leq T_{n}\leq T^{\mathrm{max}}$ for each TTD $n$), and finally $\mathbf{A}$ is the analog beamformer, typically realized through conventional phase shifters. The objective function $Q$ captures the desired QoS criterion, such as the sum spectral efficiency (sum-SE) across all carriers, and finally $G_{i},\forall i$ correspond to feasibility constraints, such as those imposed by transmit power limitations. While widely adopted for beam squint mitigation in recent works in the literature, the Sparse-TTD structure presents several challenges that are often overlooked in addressing the associated optimization problem. These limitations might significantly restrict its applicability across various practical scenarios. In the following, we outline the key practical challenges associated with this structure:
@@
  • One key limitation lies in the inherent complexity of the beamforming optimization problem. The Sparse-TTD architecture requires the joint optimization of the multi-dimensional interdependent parameter sets: the digital precoders $\mathbf{D}_{m}$ for each subcarrier $m$, the tunable delay vector $\mathbf{T}$ for the TTDs, and the analog beamformer $\mathbf{A}$. This multi-component formulation introduces significant computational complexity due to the high dimensionality and non-convex nature of the problem, particularly in large-scale systems with many carriers and antenna elements. While various approximation and convexification techniques have been proposed in the literature to address this challenge, they often result in errors due to approximations that further diverge from the ideal performance of a Full-TTD design.
  • The strong coupling among the multi-dimensional decision variables necessitates the use of decomposition strategies and multi-level alternating optimization algorithms, which are widely adopted in the literature. A common approach is to decompose the original problem into three subproblems, each optimizing one variable set while keeping the others fixed. The overall solution is then obtained via iterative alternation. While this method offers a tractable framework, it is often hindered by high computational complexity and convergence issues. Notably, most existing works do not include convergence analysis, and in many cases, there is no assurance that the iterative process converges to a stable or near-optimal solution.
@@
  • Beyond algorithmic challenges, hardware limitations for addressing the solution to the formulated optimization problem significantly influence the performance of Sparse-TTD-based hybrid beamformers. One of the most critical challenges generally overlooked in the literature is the requirement of frequency-independent wideband phase shifters corresponding to the optimal decision variables of $\mathbf{A}$. In more detail, as opposed to the digital precoders $\mathbf{D}_{m}$ where a dedicated precoder vector is assigned for each carrier $m$, the optimal value of $\mathbf{A}$ is frequency independent, which should remain fixed for all carriers across the whole bandwidth. This is a real challenging issue that can practically influence the performance and cost of the system, as elaborated in the next section.
@@ III Wideband Phase Shifters: A Bottleneck for Sparse-TTD Wideband Beamformers
As previously discussed, the optimal analog beamformer $\mathbf{A}^{*}$, corresponding to the solution of the optimization problem in (II), requires the realization of frequency-independent phase shifts that remain constant across all subcarriers, irrespective of the carrier index $m$. As the bandwidth increases, achieving such frequency-invariant phase shifts becomes increasingly challenging and costly. Depending on the fractional bandwidth, operating frequency band, and required phase resolution, various techniques may be employed to realize an approximately constant phase shift in the available wideband spectrum. These techniques generally fall into three categories: two non-TTD-based approaches and one TTD-based approach, as detailed below.
@@
The first method employs a set of wideband fixed phase shifter components to construct a dynamically tunable wide-range phase shifter, implemented using a hybrid of multiple passive architectures. A representative design is presented in [10], where a 6–18 GHz 6-bit wideband phase shifter is realized using several circuit topologies: a switched transmission line for the smallest bit (5.625), high-pass/low-pass filters for intermediate bits (11.25, 22.5, 45), a magnetically coupled all-pass network for the 90 bit, and a Lange-based structure for the 180 bit. The desired phase shift is synthesized by activating a subset of these components. The second method uses conventional narrow-band phase-shift cells, which exhibit frequency-dependent phase drift, and augments them with broadband equalizers or negative-group-delay (NGD) networks [11] leveraging circuit components having controllable impedance. These equalizer/NGD sections are designed to have a complementary group-delay-versus-frequency characteristic that compensates for the phase error across the band of interest. Finally, the third method builds upon the previous approach by replacing the narrowband phase-shift cells with wideband TTDs followed by NGD circuits [12], such that the positive group delay introduced by the TTD is compensated by the NGD.
@@
As the fractional bandwidth increases, the cost and complexity of implementing wideband phase shifters escalate significantly [13], often making the overall cost of beamformers employing the first two categories of phase shifters comparable to, or even exceeding, that of TTD-based counterparts. Furthermore, since the third method incorporates TTD units within the phase shifter structure itself, applying this approach to Sparse-TTD hybrid beamformers is fundamentally discouraged, as it increases the number of required TTDs, contradicting the primary motivation behind Sparse-TTD architectures, namely, reducing the TTD count to minimize cost and complexity. Considering the discussions presented, a key insight can be drawn regarding the practicality of Sparse-TTD beamformers:
@@
The requirement for wideband phase shifters poses a critical challenge to the cost-effectiveness and performance of Sparse-TTD hybrid beamformers in wideband and ultra-wideband scenarios, potentially offsetting their advantages over Full-TTD architectures.

### v2

@@ Abstract
Beam squint poses a fundamental challenge in wideband hybrid beamforming, particularly for mmWave and THz systems that demand both ultra-wide bandwidth and high directional beams. While conventional phase shifter-based beamformers may offer partial mitigation, True Time Delay (TTD) units provide a fundamentally more effective solution by enabling frequency-independent beam steering. However, the high cost of TTD units has recently driven much interest in Sparse-TTD architectures, which combine a limited number of TTDs with a higher number of conventional PSs to balance performance and cost. This paper provides a critical examination of beam squint mitigation strategies in wideband hybrid beamformers, comparing Full-TTD, Sparse-TTD, and Non-TTD architectures. We analyze recent Non-TTD approaches, specifically the scheme leveraging the wideband beam gain (WBBG) concept, evaluating their performance and cost characteristics against TTD-based solutions. A key focus is placed on the practical limitations of Sparse-TTD architectures, particularly the often-overlooked requirement for wideband PSs operating alongside TTDs, which can significantly impact performance and implementation cost in real-world scenarios, especially for ultra-wideband applications. Finally, we conduct a cost-performance analysis to examine the trade-offs inherent in each architecture and provide guidance on selecting the most suitable hybrid beamforming structure for various fractional bandwidth regimes.
@@ I Introduction
Efficient beamforming is essential for enabling 6G technologies that exploit mmWave and THz bands to deliver ultra-high data rates and dense connectivity.In phased array systems, analog beamforming adjusts signal phases of antenna elements utilizing a single RF chain, while digital beamforming offers fine-grained amplitude and phase control at the cost of one RF chain per antenna element. To balance performance and hardware complexity, hybrid beamforming combines both approaches by connecting fewer RF chains to large antenna arrays via analog phase shifters (PSs). This allows flexible beam control with reduced cost and power consumption. In contrast to conventional beamforming methods, which often rely on narrowband assumptions, contemporary applications increasingly demand wideband and ultra-wideband beamforming to support ultra-high data rates as well as robust communications. For instance, in an ultra-wideband multi-carrier transmitter, frequency diversity enhances robustness by allowing the system to maintain performance despite deep fading or interference on specific carriers, as unaffected carriers can ensure reliable operation [9]. This is particularly vital for frequency-selective fading channels, driving significant research into advanced wideband beamforming techniques that underpin the evolution of resilient, high-capacity 6G systems. Despite its advantages, wideband beamforming introduces a critical challenge known as beam squint. Beam squint arises when fixed phase shifts optimized for a target frequency $f_{0}$, lead to frequency-dependent beam direction change for frequency $f\neq f_{0}$. For the simple case of a uniform linear array (ULA) transmitter, the main lobe of the radiated beam at frequency $f$ is deviated by $\Delta\theta=\sin^{-1}(\frac{f_{0}}{f}\sin(\theta_{0}))-\theta_{0}$, where $\theta_{0}$ corresponds to the intended beam angle at frequency $f_{0}$ with respect to the boresight of the antenna. This frequency-dependent beam squint degrades beamforming gain and overall system performance, particularly in ultra-wideband systems.
@@
To mitigate beam squint in wideband radar and communication systems, several techniques have been proposed. The first solution scheme focuses on implementing signal processing and beamforming techniques implemented using conventional analog-domain PSs. For instance, in the context of Multi-Input Multi-Output (MIMO) wideband radar, the work in [15] addresses beam pattern design under a space-frequency nulling constraint, aiming to suppress frequency-dependent interference caused by beam squint. In [4], the authors propose a beamforming scheme for wideband radars that synthesizes frequency-invariant beam patterns via constrained waveform design, incorporating nulling and peak-to-average power ratio control in the space-frequency domain. More recently, [10] introduced a hybrid beamforming approach that broadens the beam in the angular domain to achieve max-min fair beam gain across spatially squinted beams.
@@
Wideband beamformers using conventional PSs can partially mitigate beam squint, as they cannot provide an optimal solution due to the group delay associated with such components. In this regard, the second approach aims to more fundamentally mitigate beam squint by replacing phase shifters with true time delays (TTDs). Unlike phase shifters that introduce frequency-dependent phase shifts, TTDs apply frequency-independent delays, ensuring consistent beam direction across a wide bandwidth. In general, beam squint mitigation through Non-TTD-based hybrid structures follows the general block diagram of a conventional hybrid beamformer illustrated in Fig. 1-a  [10]. This structure which employs only PSs, suffer from performance degradation when operating over large fractional bandwidths. In this regard, the second approach aims to more fundamentally mitigate beam squint by replacing PSs with true time delays (TTDs). Unlike PSs that introduce frequency-dependent phase shifts, TTDs apply frequency-independent delays, ensuring consistent beam direction across a wide bandwidth. Full-TTD architectures where each antenna element is equipped with a dedicated TTD as shown in Fig. 1-b [14], are known to optimally mitigate beam squint by providing frequency-independent delays for each antenna element. However, practical challenges such as high hardware cost makes this design inefficient for ultra-wideband applications, especially in mmWave and THz frequencies. To reduce this overhead, instead of providing all antennas with dedicated TTDs, a lower number of TTD (Sparse-TTD) can be leveraged using one of the structures illustrated in Figs. 1-c–e. The structure in Fig. 1-c places the TTD block after the RF chains and before the PSs, which allows efficient coarse delay compensation followed by per-antenna phase adjustment [13, 12]. Alternatively, in the architecture in Fig. 1-d utilized in some works such as  [7], TTDs are placed after PSs, thereby providing highly accurate delay adjustment in the final stage. This architecture, however requires a more complex structure for interconnecting the TTD block and antenna elements. Finally, Fig. 1-e places the TTDs before the RF chains  [3], allowing their operation at low-power signal levels. This reduces insertion loss and makes it feasible to employ low-power, cost-effective TTDs.
Figure 1: Wideband hybrid beamformer leveraging (a) conventional phase-shifters (PSs) without deploying TTDs (Non-TTD), (b) one TTD per antenna element (Full-TTD), and (c): sparse TTDs together with a block of PSs (Sparse-TTD)
@@
Given the relatively high cost of TTDs compared to PSs, a Sparse-TTD configuration aims to reduce the hardware cost of a Full-TTD wideband hybrid beamformer. However, this approach presents several notable drawbacks. First, it does not completely eliminate beam squint, leaving residual misalignment across some frequency bands, since full squint mitigation requires equipping each antenna with a dedicated TTD. Second, the joint optimization of precoders, TTD delays, and PSs, introduces considerable computational complexity, as the design must jointly optimize all these to achieve near-optimal performance. This is commonly followed by a series of decomposition and approximation techniques that can further degrade the performance [12, 3]. Finally, and most critically, a challenge often overlooked in the literature is the requirement for wideband PSs to operate alongside TTDs. This necessity introduces additional implementation concerns, which will be examined in detail in this work.
Following the discussions provided so far relating to inherent differences of the three aforementioned architectures, the main contributions of this paper associated with the organization visualized in Fig. 2 are summarized as follows:
@@
  • While Sparse-TTD wideband hybrid beamformers are commonly cited as a cost-effective, high-performance solution for mitigating beam squint, in this paper we critically examine the practical limitations. In particular, beyond their inherently high computational complexity, we highlight the often-overlooked requirement for wideband PSs and discuss how this constraint can significantly affect performance and implementation cost in real-world scenarios. These are explored in Sections II and III as illustrated in Fig. 2.
@@
Figure 2: The hierarchical structure of the article.
@@ II Sparse-TTD Beamforming: Implementation Challenges
$\max_{\mathbf{D}_{m},\mathbf{T},\mathbf{A}}\$$Q(\mathbf{D}_{m},\mathbf{T},\mathbf{A})$
@@
where $\mathbf{D}_{m}$ represents the digital beamformer corresponding to each carrier $m$, $\mathbf{T}$ is the vector encompassing the tunable delays of the TTDs ($0\leq T_{n}\leq T^{\mathrm{max}}$ for each TTD $n$), and finally $\mathbf{A}$ is the analog beamformer, typically realized through conventional PSs. The objective function $Q$ captures the desired QoS criterion, such as the sum spectral efficiency (sum-SE) across all carriers, and finally $G_{i},\forall i$ correspond to feasibility constraints, such as those imposed by transmit power limitations. While widely adopted for beam squint mitigation in recent works in the literature, the Sparse-TTD structure presents several challenges that are often overlooked in addressing the associated optimization problem. These limitations might significantly restrict its applicability across various practical scenarios. In the following, we outline the key practical challenges associated with this structure:
@@
$\bullet$ One key limitation lies in the inherent complexity of the beamforming optimization problem. The Sparse-TTD architecture requires the joint optimization of the interdependent parameter sets: the digital precoders $\mathbf{D}_{m}$ for each subcarrier $m$, the tunable delay vector $\mathbf{T}$ for the TTDs, and the analog beamformer $\mathbf{A}$. This multi-component formulation introduces significant computational complexity due to the high dimensionality and non-convex nature of the problem, particularly in large-scale systems with many carriers and antenna elements. While various approximation and convexification techniques have been proposed in the literature to address this challenge, they often result in errors due to approximations that further diverge from the ideal performance of a Full-TTD design.
$\bullet$ The globally optimal solution of such highly complex non-convex problems requires greedy searches over the domain of $\mathbf{A},\mathbf{T},\mathbf{D}_{m},\forall m$}, which entails exploring extremely large codebooks and is generally intractable. The strong coupling among the multi-dimensional decision variables instead motivates decomposition strategies, where multi-level alternating optimization (AO) is commonly adopted in related works (e.g., [12, 3]). In this approach, the original problem is decomposed into three subproblems, each optimizing one variable set while keeping the others fixed, and the overall solution is obtained iteratively. While this framework improves tractability, it still suffers from high computational cost and potential convergence issues. In particular, most existing studies do not provide convergence guarantees or complexity characterizations. Moreover, even if convergence can be ensured, the process is generally slow because of the large dimensionality of the three decision-variable sets and the strong coupling among them.
@@
$\bullet$ Beyond algorithmic challenges, hardware limitations for addressing the solution to the formulated optimization problem significantly influence the performance of Sparse-TTD-based hybrid beamformers. One of the most critical challenges generally overlooked in the literature is the requirement of frequency-independent wideband PSs corresponding to the optimal decision variables of $\mathbf{A}$. In more detail, as opposed to the digital precoders $\mathbf{D}_{m}$ where a dedicated precoder vector is assigned for each carrier $m$, the optimal value of $\mathbf{A}$ is frequency independent, which should remain fixed for all carriers across the whole bandwidth. This is a real challenging issue that can practically influence the performance and cost of the system, as elaborated in the next section.
@@ III Wideband Phase Shifters: A Bottleneck for Sparse-TTD Wideband Beamformers
As previously discussed, the optimal analog beamformer $\mathbf{A}^{*}$, corresponding to the solution of the optimization problem in (II), requires the realization of frequency-independent phase shifts that remain constant across all subcarriers, irrespective of the carrier index $m$. As the bandwidth increases, achieving such frequency-invariant phase shifts becomes increasingly challenging and costly. Depending on the fractional bandwidth, operating frequency band, and required phase resolution, various techniques may be employed to realize an approximately constant phase shift in the available wideband spectrum. These techniques generally fall into three categories: two non-TTD-based approaches and one TTD-based approach, as detailed below. The first method uses a set of wideband fixed PS components to construct a dynamically tunable wide-range PS, implemented using a hybrid of multiple passive architectures. A representative design is presented in [5], where a 6–18 GHz 6-bit wideband PS is realized using several circuit topologies: a switched transmission line for the smallest bit ($5.625^{\circ}$), high-pass/low-pass filters for intermediate bits ($11.25^{\circ}$, $22.5^{\circ}$, $45^{\circ}$), a magnetically coupled all-pass network for the $90^{\circ}$ bit, and a Lange-based structure for the $180^{\circ}$ bit. The second method uses conventional narrow-band phase-shift cells, which exhibit frequency-dependent phase drift, and augments them with broadband equalizers or negative-group-delay (NGD) networks [8] leveraging circuit components having controllable impedance. These equalizer/NGD sections are designed to have a complementary group-delay-versus-frequency characteristic that compensates for the phase error across the band of interest. Finally, the third method builds upon the previous approach by replacing the narrowband phase-shift cells with wideband TTDs followed by NGD circuits [11], such that the positive group delay introduced by the TTD is compensated by the NGD.
@@
As the fractional bandwidth increases, the cost and complexity of implementing wideband PSs escalate significantly [6], often making the overall cost of beamformers employing the first two categories of PSs comparable to, or even exceeding, that of TTD-based counterparts. Furthermore, since the third method incorporates TTD units within the PS structure itself, applying this approach to Sparse-TTD hybrid beamformers is fundamentally discouraged, as it increases the number of required TTDs, contradicting the primary motivation behind Sparse-TTD architectures, namely, reducing the TTD count to minimize cost and complexity. Considering the discussions presented, a key insight can be drawn regarding the practicality of Sparse-TTD beamformers:
@@
The requirement for wideband PSs poses a critical challenge to the cost-effectiveness and performance of Sparse-TTD hybrid beamformers in wideband and ultra-wideband scenarios, potentially offsetting their advantages over Full-TTD architectures.